OpenStudy (anonymous):

A vending machine offers 8 different drinks. One day, 6 employees each purchased a drink from the vending machine. Find the probability that at least 2 employees purchased the same drink. Round your answer to the nearest hundredth.

OpenStudy (ash2326):

Ok tell me fist the no. of ways 6 employees can choose 8 drinks @Emilyh117

OpenStudy (anonymous):

Each employee can get each drink @ash2326

OpenStudy (ash2326):

Suppose I'm an employee, I have 8 choices, similarly other 5 have 8 choices. Total no. of ways 6 employees can choose to have 8 drinks \[8\times 8\times 8\times 8\times 8\times 8=8^6\] Do you get this?

OpenStudy (anonymous):

Oh yeah... @ash2326

OpenStudy (anonymous):

I get it

OpenStudy (ash2326):

Now, we need to find the no. of ways, where at least two employees have same drink. So the required no. of ways= Two employees have same drink+ Three employees have same drink+four employees have same drink+5 employees have same drink+ all have same drink Do you get this?

OpenStudy (anonymous):

Yes @ash2326

OpenStudy (anonymous):

is this a combination, permutation question?

OpenStudy (ash2326):

Required no. of ways can also be found as Total ways - (no. of ways where each have a different drink) yeah this is a P and C problem

OpenStudy (anonymous):

Okay. Its a permutation problem. right?

OpenStudy (ash2326):

Both

OpenStudy (anonymous):

What???

OpenStudy (ash2326):

Both permutation and combination

OpenStudy (anonymous):

How?

OpenStudy (ash2326):

Two employees have same drink+ Three employees have same drink+four employees have same drink+5 employees have same drink+ all have same drink We need to find the no. of ways 2 employees can be selected out of 6, that's combination

OpenStudy (anonymous):

Okay. Would it be something like 8/6 8/5 8/5 8/3 8/2 8/1 ?

OpenStudy (ash2326):

Easy way is to find using this At least Two employees have same drink=Total ways - (no. of ways where each have a different drink)

OpenStudy (anonymous):

@ash2326

OpenStudy (ash2326):

no. of ways where each have a different drink First person has 8 choices second has 7 third has 6 and so on No. of ways where each have a different drink=8*7*6*5*4*3

OpenStudy (ash2326):

Do you get it?

OpenStudy (anonymous):

I think so @ash2326